OpenStudy (anonymous):

is it possible to get a decimal for the problem: y^2=5^2 + [5(radical)9]^2

OpenStudy (valpey):

You can get precision to any number of digits, but this is still irrational. \[y^2 = 5^2 +(5\sqrt{9})^2 = 25+15^2=250 = 25*10\] \[y = \pm 5\sqrt{10}\] Can you derive a decimal representation of the square root of 10?

OpenStudy (anonymous):

im sorry i didnt get that

OpenStudy (anonymous):

\[y ^{2}=5^{2}+(5\sqrt{3})^{2}\]